# DisenSemi: Semi-supervised Graph Classification via Disentangled Representation Learning

### v1

@@ Abstract
Graph classification is a critical task in numerous multimedia applications, where graphs are employed to represent diverse types of multimedia data, including images, videos, and social networks. Nevertheless, in real-world scenarios, labeled graph data can be limited or scarce. To address this issue, we focus on the problem of semi-supervised graph classification, which involves both supervised and unsupervised models learning from labeled and unlabeled data. In contrast to recent approaches that transfer the entire knowledge from the unsupervised model to the supervised one, we argue that an effective transfer should only retain the relevant semantics that align well with the supervised task. In this paper, we propose a novel framework named DisenSemi, which learns disentangled representation for semi-supervised graph classification. Specifically, a disentangled graph encoder is proposed to generate factor-wise graph representations for both supervised and unsupervised models. Then we train two models via supervised objective and mutual information (MI)-based constraints respectively. To ensure the meaningful transfer of knowledge from the unsupervised encoder to the supervised one, we further define an MI-based disentangled consistency regularization between two models and identify the corresponding rationale that aligns well with the current graph classification task. Experimental results on a range of publicly accessible datasets reveal the effectiveness of our DisenSemi.
@@ I Introduction
Graphs are among the most versatile data structures in many real-world multimedia applications across various domains, e.g., biology networks [1], molecule structures [2], citation networks [3], social networks [4] and recommender system [5], etc. For graph-structured data, one fundamental task is graph classification, which endeavors to capture the property of the entire graph and has become an important research hot spot in numerous multimedia tasks, including chemical compound prediction [2], text categorization [6] and social network analysis [4].
@@
To tackle the graph classification problem, many machine learning-based methods are proposed. Graph kernel methods leverage structural patterns such as shortest paths [7], subtrees [8], subgraphs [9] and graphlets [10] to measure the similarity among graphs and pass the similarity matrix to a kernel-based supervised algorithm such as the Support Vector Machine (SVM) to perform classification. Instead of hand-crafted feature extraction in graph kernel methods, more recently, graph neural networks (GNNs) can extract graph structural features in a supervised end-to-end manner. Despite the better performance, GNNs always require a large amount of labeled data, which is highly expensive and time-consuming in the real world [11]. For example, as demonstrated in Fig. 1, the label of a chemical compound is typically produced with a costly Density Functional Theory (DFT) calculation or several complicated experiments.
Figure 1: Illustration of the molecular complicated annotating procedure. It takes around four to five hours to compute the characteristics of the molecule, which fully demonstrates the graph label scarcity in reality.
@@
In practice, there always exist massive unlabeled graph samples. Although labels of these graphs are unavailable, the inherent knowledge of the graph may help to enhance GNNs’ encoding more expressive and discriminative. Semi-supervised classification methods combine both supervised and unsupervised models where the unsupervised model learns from unlabeled data and can serve as a regularizer. Indeed, there are a handful of semi-supervised works proposed for graph classification, which can be divided into several categories, i.e., pseudo-labeling [12, 13, 14], consistency learning [15, 16, 17, 18, 2, 19, 20]. Especially for consistency learning, methods like InfoGraph [16] and GraphSpa [20] jointly learn supervised and unsupervised models from labeled and unlabeled data. Meanwhile, a popular fashion (i.e., GraphCL [15] and JOAO [17]) has been developed recently which first pre-trains the unsupervised model with unlabeled data and then fine-tunes the model with labeled data.
@@
However, these existing semi-supervised graph classification methods still remain unsatisfactory due to the following limitations. (1) Neglect the entanglement of latent factors. The formation of graphs in the real world typically stems from the highly complex interaction of different latent factors. For example, a molecular graph may consist of various groups of atoms and bonds representing different functional units [21]. Existing works often neglect the entanglement of latent factors and the extracted features are holistic, which harms interpretability and leads to sub-optimal performance for the graph classification task. (2) Mismatch semantics between supervised and unsupervised tasks. Supervised and unsupervised tasks on labeled and unlabeled graph data may capture different information or occupy separate semantic spaces, corresponding to distinct graph factors. Simply combining two tasks may lead to a “negative transfer” [22]. Therefore, we argue that a more efficient way is to disentangle the graph representation learned from unsupervised tasks into distinct latent factors and transfer the corresponding factor-wise information well aligned to semantically enhance the supervised task for semi-supervised graph classification.
@@
Toward this end, in this paper, we propose DisenSemi, a novel Disentangled graph representation learning framework for Semi-supervised graph classification, which is capable of enhancing the supervised prediction task with the unsupervised representation learning tasks semantically. In contrast to existing works that directly utilize an entangled representation learned from a large amount of unlabeled data for prediction, our proposed approach emphasizes regularizing the rationale to explicitly exploit transferable factor information between supervised and unsupervised tasks. In particular, we design a disentangled graph encoder by characterizing the global-level topological semantics, during which the whole graph is first decomposed into factor graphs. Then the factor-wise interpretable graph representation is produced via the multi-channel message-passing layer, where each channel is tailored to propagate features under one factor graph and a separate readout operation in each channel summarizes the specific aspect of the graph. For labeled data, we train the model using a supervised objective function while for unlabeled data, the model is trained with MI-based constraints for the input factorized graph and its corresponding features to ensure the disentanglement. Next we conduct MI maximization between the supervised and unsupervised models under each latent factor instead of in the whole feature space for disentangled consistency regularization. Compared with the existing works, this novel factor-wise MI estimation strategy can ensure the regularized factor is best pertinent to the aspect bridging supervised and unsupervised models for the current semi-supervised graph classification task. More importantly, we demonstrate that our framework can be formalized as a problem of maximizing the log-likelihood solved by Expectation Maximization (EM).
To summarize, in this paper we make the following contributions:
  • Conceptual: We propose a novel disentangled representation learning framework for semi-supervised graph classification, which explicitly models the rationale factor that fits well between supervised and unsupervised learning tasks. And the proposed mechanism can be potentially generalized to other semi-supervised learning tasks.
@@
  • Experimental: We conduct extensive experiments on a variety of public datasets to evaluate the DisenSemi. Experimental results demonstrate the efficiency and outstanding interpretability of our proposed framework for semi-supervised graph classification tasks.
@@ II-A Semi-supervised Graph Classification
Semi-supervised learning has received growing attention in recent years. It is associated with a learning paradigm that uses both labeled and unlabeled data and encompasses two prominent techniques. The first line is the consistency regularization methods, which are based on the manifold or smoothness assumption and posit that realistic perturbations of the data points should not change the output of the model. The most common structure is Teacher-Student, consisting of two models called student and teacher and applying a consistency constraint between the two predictions of student and teacher [29, 30, 31]. The second line is pseudo-labeling methods, which predict the label distribution of unlabeled samples and select confident samples to provide further guidance during the training process [32, 33].
For graph-structured data, there are also attempts using consistency regularization [16, 15, 18, 14], pseudo-labeling [34, 13, 12, 35] or other approaches [36, 37, 38, 39, 20] for semi-supervised learning. Typically, InfoGraph [16] learns supervised and unsupervised model respectively and estimates the MI between two models. GraphCL [15] and GLA [14] use contrastive learning to get graph representations for semi-supervised graph classification. DualGraph [13] and UGNN [35] jointly learn a prediction and a retrieval module via posterior regularization during the pseudo-labeling process. GraphSpa [20] actively selects informative graphs for subsequent model training using a hybrid selection strategy. However, these approaches fail to disentangle the diverse underlying factors behind the graph data, making it challenging to identify suitable alignments between supervised and unsupervised models.
@@ II-B Disentangled Representation Learning
Disentangled representation learning aims to acquire factorized representations capable of discerning and separating the latent explanatory factors concealed within the observed data [23]. By achieving such a disentangled representation, the model gains a deeper understanding of the intricate dependencies within the data, facilitating more insightful and meaningful analysis or manipulation of the underlying factors. Existing efforts about disentangled representation learning are mainly on computer vision [40, 41], natural language processing [42, 3] and recommendation [24, 43, 44].
Recently, there has been a notable surge of interest in applying disentangled representation learning techniques to graph-structured data [25, 5, 26, 27]. DisenGCN [25] learns disentangled node representation through a neighborhood routing mechanism to divide the neighborhood of the node into several mutually exclusive parts. DisenHAN [5] learns disentangled representation in Heterogeneous Information Network (HIN) to iteratively identify the major aspect of the relation between node pairs and propagate corresponding information semantically. FactorGCN [26] takes the whole graph as input and produces block-wise interpretable graph-level features for classification. DGCL [27] proposes a self-supervised graph disentangled representation framework via contrastive learning method. UMGRL [28] provides a self-supervised disentangled representation learning method for the multiplex graph to capture common and private graph information. Our work focuses on learning disentangled representation for semi-supervised graph classification. And our primary objective is to design a framework where a supervised model can learn from an unsupervised model through the latent semantic space that aligns well with the current graph classification task.
@@ III-A Problem Definition
Definition 1 (Graph). A graph is formally defined as $G=(V,E,X)$, where $V$ and $E\in V\times V$ are the node set and edge set of the graph respectively. We denote the node feature matrix as $X\in\mathbb{R}^{|V|\times d^{\prime}}$, where each row $x_{i}\in\mathbb{R}^{d^{\prime}}$ represents the feature vector of node $i$ and $d^{\prime}$ is the dimension of node feature. A graph is typically labeled if it is associated with a class label $y\in\{0,1\}^{C}$, where $C$ is the number of classes. A graph is unlabeled if its class label is unknown.
Definition 2 (Semi-supervised Graph Classification). Given a set of graphs $\mathcal{G}=\{\mathcal{G}^{L},\mathcal{G}^{U}\}$, where $\mathcal{G}^{L}=\{G_{1},\dots,G_{|\mathcal{G}^{L}|}\}$ and $\mathcal{G}^{U}=\{G_{|\mathcal{G}^{L}|+1},\dots,G_{|\mathcal{G}^{L}|+|\mathcal{G}^{U}|}\}$ represent labeled and unlabeled graphs respectively. Let $\mathcal{Y}^{L}=\{y_{1},\dots,y_{|\mathcal{G}^{L}|}\}$ represents the label corresponding to $\mathcal{G}^{L}$. Semi-supervised graph classification seeks to learn a prediction function that can assign class labels to unlabeled graph data in $\mathcal{G}^{U}$ based on the available class labels in $\mathcal{G}^{L}$.
@@ III-B Preliminaries
Preliminary 1 (MI Estimation). MI is a fundamental measurement of the dependence between two random variables, which has been applied to a wide range of tasks [41, 45]. However, estimation MI is intractable especially in high dimensional continuous settings. Since earlier non-parametric binning methods perform poor, a neural estimator MINE [46] formates MI between $x$ and $y$ as Kullback-Leibler (KL)-divergence between the joint distribution $p_{xy}$ and the product of the marginals $p_{x}\otimes p_{y}$,
@@
and derive a lower bound of MI with dual representations of KL-divergence. Following previous works [45, 16], we can further rely on non-KL divergence (i.e., Jensen-Shannon divergences) as an alternative for the MI estimator.
@@
where $T(\cdot,\cdot)$ is a critic (or score) function approximated by a neural network. In this way, the loss can be seen as a standard binary cross-entropy (BCE) loss between the samples from the joint (a.k.a. positive examples) and the product of marginals (a.k.a. negative examples).
Preliminary 2 (MI and Contrastive Learning). More recently, InfoNCE [47] propose a Noise Contrastive Estimation (NCE [48])-based MI lower bound, defined as:
$I_{NCE}(x,y):=\mathbb{E}\Bigl{[}\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp^{T(x_{i},y_{i})}}{\frac{1}{N}\sum_{j=1}^{N}\exp^{T(x_{j},y_{j})}}\Bigr{]},$ (3)
where the expectation is over $N$ samples $\{(x_{i},y_{i})\}_{i=1}^{N}$ drawn from the joint distribution $p_{xy}$. And contrastive learning encourages learning the model where representations between positive samples are pulled closer (e.g., graph representations corresponding to the same factor bridging supervised and unsupervised tasks) and representations among negative samples are pushed apart.
@@ IV-A Overview
where $L_{S}$ denotes the supervised loss that measures the discrepancy between the supervised model prediction and the real label of the graph data, $L_{U}$ denotes the unsupervised loss acting as a regularization term and $\lambda$ is the relative weight between two losses.
However, we argue that supervised and unsupervised tasks have different optimization targets, which correspond to different semantic spaces of the graph data. Thus, we propose DisenSemi for semi-supervised graph classification. The basic idea is to explicitly infer the latent factors underlying a large amount of unlabeled graph data and transfer the well-aligned knowledge to enhance the supervised task in a factor-wise manner. As shown in Fig. 2, our framework is composed of a supervised model and an unsupervised model, where graph-level representations of both models are learned via the GNN-based encoder. Given the graph data, the GNN-based encoder first factorizes the input graph into several factor graphs, then gets graph representation by propagating and aggregating features on the corresponding factor graph. For the supervised model, we merge all the extracted factor graph representations to predict graph labels. For the unsupervised model, we introduce several MI-based constraints among factorized graphs and their corresponding extracted representations. Finally, a disentangled consistency regularization is conducted to explicitly identify the rationale between the two models and transfer the learned knowledge semantically via factor-wise MI maximization. The process is equivalent to a Variational EM algorithm maximizing the log-likelihood.
Figure 2: A schematic view of the DisenSemi, which consists of a supervised and an unsupervised model. We extract factor-wise graph representations of two models and maximize the agreement under each latent factor.

### v2

@@ Abstract
Graph classification is a critical task in numerous multimedia applications, where graphs are employed to represent diverse types of multimedia data, including images, videos, and social networks. Nevertheless, in the real world, labeled graph data is always limited or scarce. To address this issue, we focus on the semi-supervised graph classification task, which involves both supervised and unsupervised models learning from labeled and unlabeled data. In contrast to recent approaches that transfer the entire knowledge from the unsupervised model to the supervised one, we argue that an effective transfer should only retain the relevant semantics that align well with the supervised task. We introduce a novel framework termed DisenSemi in this paper, which learns disentangled representation for semi-supervised graph classification. Specifically, a disentangled graph encoder is proposed to generate factor-wise graph representations for both supervised and unsupervised models. Then we train two models via supervised objective and mutual information (MI)-based constraints respectively. To ensure the meaningful transfer of knowledge from the unsupervised encoder to the supervised one, we further define an MI-based disentangled consistency regularization between two models and identify the corresponding rationale that aligns well with the current graph classification task. Experiments conducted on various publicly available datasets demonstrate the effectiveness of our DisenSemi.
@@ I Introduction
Graphs are among the most versatile data structures in many real-world multimedia applications across various domains, e.g., biology networks [1], molecule structures [2], citation networks [3], social networks [4] and recommender system [5], etc. One fundamental task for graph-structured data is graph classification, which endeavors to capture the characteristics of the entire graph and has become an important research hot spot in numerous multimedia tasks, including chemical compound prediction [2], text categorization [6] and social network analysis [4].
@@
To tackle the graph classification problem, many machine learning-based methods are proposed. Graph kernel methods leverage structural patterns such as shortest paths [7], subtrees [8], subgraphs [9] and graphlets [10] to measure similarity among graphs. These methods then use the similarity matrix with a kernel-based supervised algorithm, like the Support Vector Machine (SVM), to perform classification. Instead of hand-crafted feature extraction in graph kernel methods, more recently, graph neural networks (GNNs) can extract graph structural features in a supervised end-to-end manner. Despite the superior performance, GNNs often require a substantial amount of labeled data, which can be costly and time-consuming in real-world scenarios [11].
@@
In practice, there always exist massive unlabeled graph samples. Although labels of these graphs are unavailable, the inherent knowledge of the graph may help to enhance GNNs’ encoding more expressive and discriminative. Semi-supervised classification methods combine both supervised and unsupervised models where the unsupervised model learns from unlabeled data and can serve as a regularizer. Indeed, there are a handful of semi-supervised works proposed for graph classification, which can be categorized into several types, i.e., pseudo-labeling [12, 13, 14], consistency learning [15, 16, 17, 18, 2, 19]. Especially for consistency learning, methods like InfoGraph [16] and GraphSpa [19] jointly learn supervised and unsupervised models from labeled and unlabeled data. Meanwhile, a popular fashion (i.e., GraphCL [15] and JOAO [17]) has been developed recently, which first pre-train the model in an unsupervised manner using unlabeled data and then fine-tune the model with labeled data.
@@
However, these previous semi-supervised graph classification methods still face the following limitations. (1) Neglect the intricate interactions of latent factors. In the real world, the formation of graphs often results from the intricate interactions of various latent factors. For instance, a molecular graph may include diverse groups of atoms and bonds that represent distinct functional units [20]. Existing works often overlook the interactions of latent factors and the extracted features are holistic, which harms interpretability and leads to sub-optimal performance for the graph classification task. (2) Mismatch semantics between supervised and unsupervised tasks. Supervised and unsupervised tasks on labeled and unlabeled graph data may capture different information or occupy separate semantic spaces, corresponding to distinct graph factors. Simply combining two tasks may lead to a “negative transfer” [21]. Therefore, we argue that a more efficient way is to disentangle the graph representation learned from unsupervised tasks into distinct latent factors and transfer the corresponding factor-wise information well aligned to semantically enhance the supervised task for semi-supervised graph classification.
@@
Toward this end, we propose DisenSemi in this paper, a novel Disentangled graph representation learning framework for Semi-supervised graph classification. This framework enhances the supervised prediction task with the unsupervised representation learning tasks in a semantically meaningful way. In contrast to existing works that directly utilize an entangled representation learned from unlabeled data for prediction, our proposed approach emphasizes regularizing the rationale to explicitly exploit transferable factor information between supervised and unsupervised tasks. Specifically, we propose a disentangled graph encoder that characterizes global-level topological semantics. The encoder first decomposes the entire graph into several factor graphs. Then the factor-wise interpretable graph representation is extracted through multiple channels of the message-passing layer. Each channel is specifically designed to propagate features within one factor graph, with a separate readout operation in each channel summarizing a distinct aspect of the graph. For labeled data, we train the model using a supervised objective function while for unlabeled data, the model is trained with MI-based constraints for the input factorized graph and its corresponding features to ensure the disentanglement. Next we conduct MI maximization between the supervised and unsupervised models under each latent factor instead of in the whole feature space for disentangled consistency regularization. Compared with the existing works, this novel factor-wise MI estimation strategy can ensure the regularized factor is best pertinent to the aspect bridging supervised and unsupervised models for the current semi-supervised graph classification task. More importantly, we demonstrate that our framework can be formalized as a problem of maximizing the log-likelihood solved by Expectation Maximization (EM).
To summarize, we make the following contributions:
  • Conceptual: We propose a novel disentangled representation learning framework for the semi-supervised graph classification task, which explicitly models the rationale factor that fits well between supervised and unsupervised learning models. And the proposed framework can be applied to other semi-supervised learning tasks.
@@
  • Experimental: We perform extensive experiments on a range of public datasets to evaluate the performance of DisenSemi. Experimental results demonstrate the efficiency and outstanding interpretability of our proposed framework for semi-supervised graph classification tasks.
@@ II-A Semi-supervised Graph Classification
Semi-supervised learning has garnered increasing attention in recent years. It is associated with a paradigm that learns from both labeled and unlabeled data and encompasses two prominent techniques. The first line is the consistency regularization methods, which rely on the manifold or smoothness assumption and posit that minor perturbations of the data points should not affect the model’s output. The most common structure is the Teacher-Student framework, which involves two models: the student and the teacher. This framework applies a consistency constraint between the predictions made by the student and the teacher models [28, 29, 30]. The second line is pseudo-labeling methods, which involve predicting the label distribution for unlabeled samples and selecting the most confident samples to provide additional guidance during the training process [31, 32].
For graph-structured data, there are also attempts using consistency regularization [16, 15, 18, 14, 11], pseudo-labeling [33, 13, 12, 34] or other approaches [35, 19] for semi-supervised learning. Typically, InfoGraph [16] learns supervised and unsupervised model respectively and estimates the MI between two models. GraphCL [15] and GLA [14] use contrastive learning to get graph representations, which are then used in the fine-tuning step for semi-supervised graph classification. DualGraph [13] and UGNN [34] jointly learn the prediction and retrieval modules via posterior regularization during the pseudo-labeling process. GraphSpa [19] employs an active learning approach to select informative graphs for semi-supervised model training. However, these approaches fail to disentangle the diverse underlying factors behind the graph data, making it challenging to identify suitable alignments between supervised and unsupervised models.
@@ II-B Disentangled Representation Learning
The goal of disentangled representation learning is to acquire factorized representations capable of discerning and separating underlying latent factors within the observed data [22]. Existing efforts in disentangled representation learning are primarily focused on computer vision [36, 37], natural language processing [38, 3] and recommendation [23, 39].
Recently, there has been a notable surge of interest in applying disentangled representation learning techniques to graph-structured data [24, 5, 25, 26]. DisenGCN [24] learns disentangled node representations using a neighborhood routing mechanism, which partitions the node’s neighborhood into several mutually exclusive parts. DisenHAN [5] learns disentangled representation in Heterogeneous Information Network (HIN) by iteratively identifying the primary aspects of the relationships between node pairs and semantically propagating the corresponding information. FactorGCN [25] takes the whole graph as input and produces block-wise interpretable graph-level features for classification. DGCL [26] proposes a self-supervised graph disentangled representation framework via the contrastive learning method. UMGRL [27] provides a self-supervised disentangled representation learning method for the multiplex graph to capture common and private graph information. Our work focuses on learning disentangled representation for the semi-supervised graph classification task. And our primary objective is to design a framework where a supervised model can learn from an unsupervised model through the latent semantic space that aligns well with the current graph classification task.
@@ III-A Problem Definition
Definition 1 (Graph). We define a graph as $G=(V,E,X)$, where $V$ and $E\in V\times V$ denote the node set and edge set of the graph, respectively. $X\in\mathbb{R}^{|V|\times d^{\prime}}$ denote the node feature matrix, where each row $x_{i}\in\mathbb{R}^{d^{\prime}}$ represents the initial feature of node $i$ and $d^{\prime}$ is the node feature dimension. A graph is typically labeled if it contains a class label $y\in\{0,1\}^{C}$, where $C$ denotes the number of classes. Otherwise, the graph is considered unlabeled when the class label is unknown.
Definition 2 (Semi-supervised Graph Classification). Given a set of graphs $\mathcal{G}=\{\mathcal{G}^{L},\mathcal{G}^{U}\}$, where $\mathcal{G}^{L}=\{G_{1},\dots,G_{|\mathcal{G}^{L}|}\}$ and $\mathcal{G}^{U}=\{G_{|\mathcal{G}^{L}|+1},\dots,G_{|\mathcal{G}^{L}|+|\mathcal{G}^{U}|}\}$ represent labeled and unlabeled graphs respectively. Let $\mathcal{Y}^{L}=\{y_{1},\dots,y_{|\mathcal{G}^{L}|}\}$ represents the label corresponding to $\mathcal{G}^{L}$. Semi-supervised graph classification seeks to learn a prediction function that can assign class labels to the unlabeled graph data in $\mathcal{G}^{U}$ based on the class labels available in $\mathcal{G}^{L}$.
@@ III-B Preliminaries
Preliminary 1 (MI Estimation). MI is a key metric for assessing the dependence between two random variables, and it has been utilized across a diverse array of tasks [37, 40]. However, estimation MI is intractable especially in high dimensional continuous settings. Since earlier non-parametric binning methods perform poor, a neural estimator MINE [41] formates MI between $x$ and $y$ as Kullback-Leibler (KL)-divergence between their joint distribution $p_{xy}$ and the product of their marginal distributions $p_{x}\otimes p_{y}$,
@@
and derive a lower bound of the KL-divergence for estimation. Following previous works [40, 16], we can further rely on non-KL divergence (i.e., Jensen-Shannon divergences) as an alternative for the MI estimator.
@@
where $T(\cdot,\cdot)$ is a critic (or score) function approximated by a neural network. In this way, the loss can be seen as a standard binary cross-entropy (BCE) loss between positive samples from the joint distribution and negative samples from the product of the marginals.
Preliminary 2 (MI and Contrastive Learning). More recently, InfoNCE [42] propose a Noise Contrastive Estimation (NCE [43])-based lower bound for MI estimation, defined as:
$I_{NCE}(x,y):=\mathbb{E}\Bigl{[}\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp^{T(x_{i},y_{i})}}{\frac{1}{N}\sum_{j=1}^{N}\exp^{T(x_{i},y_{j})}}\Bigr{]},$ (3)
where $N$ samples $\{(x_{i},y_{i})\}_{i=1}^{N}$ drawn from $p_{xy}$ are used to compute the expectation. Contrastive learning encourages the model where representations between positive samples are pulled closer (e.g., graph representations corresponding to the same factor bridging supervised and unsupervised tasks) and representations among negative samples are pushed apart.
@@ IV-A Overview
where $L_{S}$ denotes the supervised loss, which quantifies the difference between the predictions of the supervised model and the actual labels of the graph data, $L_{U}$ denotes the unsupervised loss acting as a regularization term and $\lambda$ is the relative weight between two losses.
However, we argue that supervised and unsupervised tasks have different optimization targets, which correspond to different semantic spaces of the graph data. Thus, we propose DisenSemi for semi-supervised graph classification. The basic idea is to explicitly infer the latent factors underlying a substantial amount of unlabeled graph data and transfer the well-aligned knowledge to enhance the supervised task in a factor-wise manner. As shown in Fig. 1, our framework is composed of a supervised model and an unsupervised model, where graph-level representations of both models are learned via the GNN-based encoder. Given the graph data, the GNN-based encoder first factorizes the input graph into several factor graphs, then gets graph representation by propagating and aggregating features on the corresponding factor graph. For the supervised model, we merge all the extracted factor graph representations to predict graph labels. For the unsupervised model, we introduce several MI-based constraints among factorized graphs and their corresponding extracted representations. Finally, a disentangled consistency regularization is conducted to explicitly identify the rationale between the two models and transfer the learned knowledge semantically via factor-wise MI maximization. The process is equivalent to a Variational EM algorithm maximizing the log-likelihood.
Figure 1: A schematic view of the DisenSemi, which consists of a supervised and an unsupervised model. We extract factor-wise graph representations from two models and maximize the agreement in a factor-wise manner.